OpenStudy (anonymous):

A motorcyclist who is moving along an x axis directed toward the east has an acceleration given by a = (6.30 - 2.20t) m/s2 for 0 ≤ t ≤ 6.0 s. At t = 0, the velocity and position of the cyclist are 3.00 m/s and 7.3 m. What total distance does the cyclist travel between t = 0 and 6.0 s? idk why but i keep getting the wrong answer.

OpenStudy (anonymous):

What answer do you come up with?

OpenStudy (anonymous):

51.5

OpenStudy (michele_laino):

we have to integrate the acceleration first, in order to get the velosity function: \[v\left( t \right) = \int {a\left( t \right)dt} = \int {\left( {6.3 - 2.2t} \right)dt} = ...?\]

OpenStudy (michele_laino):

oops... the velocity function...

OpenStudy (anonymous):

so it would be 6.3t - (2.2/2)t^2 ?

OpenStudy (michele_laino):

we have to add the arbitrary constant, namely: \[v\left( t \right) = 6.3t - 1.1{t^2} + C\]

OpenStudy (michele_laino):

next, we have to apply the initial condition, nameli v(0) = 3. So we get: \[\begin{gathered} 3 = v\left( 0 \right) = 6.3 \cdot 0 - 1.1 \cdot 0 + C = C \hfill \\ C = 3 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

finally our velocity function will be this: \[v\left( t \right) = 6.3t - 1.1{t^2} + 3\]

OpenStudy (anonymous):

and then put t = 6 in the equation?

OpenStudy (anonymous):

so the answer is 1.2?

OpenStudy (michele_laino):

no, we have to integrate the velocity function, in order to get the position function x(t), namely: \[x\left( t \right) = \int {\left( {6.3t - 1.1{t^2} + 3} \right)dt} \]

OpenStudy (anonymous):

oh the answer is 37.2?

OpenStudy (michele_laino):

you should get this: \[x\left( t \right) = \int {\left( {6.3t - 1.1{t^2} + 3} \right)dt} = \frac{{6.3}}{2}{t^2} - \frac{{1.1}}{3}{t^3} + 3t + {C_1}\]

OpenStudy (michele_laino):

where C_1 is another arbitrary constant, whose value is given applying the initial condition for the position x, namely: \[\begin{gathered} 7.3 = x\left( 0 \right) = \frac{{6.3}}{2} \cdot 0 - \frac{{1.1}}{3} \cdot 0 + 3 \cdot 0 + {C_1} = {C_1} \hfill \\ {C_1} = 7.3 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

finally, the position function, will be this: \[x\left( t \right) = \frac{{6.3}}{2}{t^2} - \frac{{1.1}}{3}{t^3} + 3t + 7.3\]

OpenStudy (michele_laino):

The requested distance is therefore: \[\begin{gathered} x\left( 6 \right) - x\left( 0 \right) = \hfill \\ = \left[ {\frac{{6.3}}{2} \cdot {6^2} - \frac{{1.1}}{3} \cdot {6^3} + 3 \cdot 6 + 7.3} \right] - \left[ {\frac{{6.3}}{2} \cdot 0 - \frac{{1.1}}{3} \cdot 0 + 3 \cdot 0 + 7.3} \right] = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

-675?

OpenStudy (michele_laino):

I got 52.2, is it right?

OpenStudy (anonymous):

yes! thank you :)

OpenStudy (michele_laino):

thank you!